OpenStudy (anonymous):

How many grams of methane gas (CH4) are in a 15.8 liter sample at 1.2 atmospheres and 27°C? Show all work used to solve this problem

OpenStudy (somy):

PV=nRT

OpenStudy (somy):

can you use it?

OpenStudy (somy):

on your own i mean

OpenStudy (anonymous):

I have no idea what that means

OpenStudy (somy):

its a formula for ideal gases

OpenStudy (somy):

PV= nRT p- pressure V- volume n- mole R- constant T- temperature

OpenStudy (somy):

so what are you given in the question ?

OpenStudy (anonymous):

I don't know, I know we have 15.8 liters

OpenStudy (somy):

so ur volume 15.8 L

OpenStudy (somy):

pressure is also given

OpenStudy (somy):

and temperature

OpenStudy (somy):

so P= 1.2 atm V= 15.8 L T= 27 °C

OpenStudy (somy):

you need to change units for pressure change it from atm to Pa for volume change it from L to m3 for temperature change it from °C to Kelvin

OpenStudy (somy):

can u do that?

OpenStudy (anonymous):

If you tell me how haha

OpenStudy (somy):

lol okay

OpenStudy (somy):

atm to pascal you need to multiply by 101 325

OpenStudy (somy):

L to m3 you need to divide by 1000

OpenStudy (somy):

and °C to Kelvin you need to add 273

OpenStudy (somy):

now do it and tell me the new values

OpenStudy (anonymous):

Where did you get these numbers?

OpenStudy (somy):

this is conversion

OpenStudy (somy):

you need to learnt how to change units

OpenStudy (anonymous):

So the new temperature is 300?

OpenStudy (somy):

yes :D

OpenStudy (anonymous):

and the new volume is 0.0158?

OpenStudy (somy):

thats right :)

OpenStudy (anonymous):

And I don't understand the atm one, you have two numbers next to each other

OpenStudy (somy):

come again? what do u mean 2 numbers next to each other?

OpenStudy (anonymous):

How do I convert atm to whatever I'm supposed to convert it to

OpenStudy (somy):

what is the value for atm?

OpenStudy (anonymous):

1.2

OpenStudy (somy):

good so if 1 atm = 101 325 1.2 atm= ?

OpenStudy (anonymous):

I don't know what 101 325 means are you missing a comma or something

OpenStudy (somy):

1 atm = 101 325 Pa its number...

OpenStudy (gazela1997):

101.325 KPa Change the Celcius to kelivin

OpenStudy (somy):

she did.

OpenStudy (somy):

so it'll be 1.2*101 325 =

OpenStudy (anonymous):

that's so confusing, 121590?

OpenStudy (somy):

yes

OpenStudy (somy):

so now your new values P= 121590 Pa V= 0.0158 m3 T= 300 K

OpenStudy (somy):

right?

OpenStudy (anonymous):

Right

OpenStudy (somy):

you have to know that R is constant and the value of R is 8.31

OpenStudy (somy):

get it?

OpenStudy (anonymous):

Okay

OpenStudy (somy):

cool so the formula we'll be using is PV=nRT what you don't have?

OpenStudy (anonymous):

So basically pressure x volume = R x temp?

OpenStudy (somy):

yes but don't forget 'n'

OpenStudy (somy):

pressure*volume= mole*R*temperature

OpenStudy (anonymous):

n means mole?

OpenStudy (somy):

yes

OpenStudy (somy):

so we don't have mole right?

OpenStudy (anonymous):

Okay so pressure*volume = 18.96?

OpenStudy (somy):

okay first can u make mole as subject?

OpenStudy (anonymous):

I don't know what that means

OpenStudy (somy):

for example if i want to find V , i'll make it as subject and the formula will look like this V= nRT/P

OpenStudy (anonymous):

So I'm looking for n = something?

OpenStudy (somy):

yes

OpenStudy (somy):

so if u make n as subject how the formula will look like?

OpenStudy (anonymous):

n = R*temperature ?

OpenStudy (somy):

how about volume and pressure?

OpenStudy (anonymous):

What about it?

OpenStudy (somy):

your main formula is PV= nRT

OpenStudy (anonymous):

Okay

OpenStudy (somy):

so if u make n as subject it'll be like this n= PV/RT

OpenStudy (somy):

do u understand?

OpenStudy (anonymous):

Okay

OpenStudy (somy):

now put the values into the formula and get your mole

OpenStudy (anonymous):

So n = 1921. 122/2493 = 0.770 ?

OpenStudy (somy):

yup :) 0.77

OpenStudy (somy):

good job :)

OpenStudy (somy):

now u found mole

OpenStudy (somy):

can you tell me what is the formula of mole?

OpenStudy (anonymous):

Mass/molecular mass?

OpenStudy (somy):

good :)

OpenStudy (somy):

what is the question asking you to find?

OpenStudy (anonymous):

how many grams of CH4 are in a 15.8 liter sample at 1.2 atmospheres and 27°C?

OpenStudy (somy):

so they are asking for Mas right?

OpenStudy (somy):

mass*

OpenStudy (anonymous):

Yeah they are

OpenStudy (somy):

mole= mass/Mr make mass as subject

OpenStudy (anonymous):

I'm not good at that haha but mass = Mr/n?

OpenStudy (somy):

no mass= Mr*mole

OpenStudy (somy):

now what is Mr of CH4?

OpenStudy (anonymous):

That was my next guess, really quick Mr is what?

OpenStudy (somy):

Mr is molecular mass

OpenStudy (anonymous):

Okay where do I find it is it on the periodic table?

OpenStudy (somy):

yes

OpenStudy (anonymous):

16.04 g/mol

OpenStudy (somy):

yeah basically 16

OpenStudy (anonymous):

Okay so mass = 16 x 0.770 = 12.32

OpenStudy (somy):

yup :D

OpenStudy (somy):

done :)

OpenStudy (anonymous):

Now what or are we done?

OpenStudy (anonymous):

Yay! Do you mind if i write my answer in paragragh form and than I'll tag you in it and you can tell me if it looks good?

OpenStudy (somy):

okay np :)